# New Factorizations of Yang-Mills Amplitudes

### v1

@@ Abstract
We propose a new factorization pattern for tree-level Yang-Mills (YM) amplitudes, where they decompose into a sum of products of two lower-point amplitudes by setting specific two-point non-planar Mandelstam variables within a rectangular configuration to zero. This approach manifests the hidden zeros of YM amplitudes recently identified. Furthermore, by setting specific Lorentz products involving polarization vectors to zero, the amplitudes further reduce to a sum of products of three currents. These novel factorizations provide a fresh perspective on the structure of YM amplitudes, potentially enhancing our understanding and calculation of these fundamental quantities.
@@ Introduction
At the loop level, (generalized) unitarity cuts are employed to determine the loop integrand of amplitudes Bern et al. (1994, 1995). The primary difference from conventional factorization is the substitution of a loop propagator for a tree propagator on-shell. Thus, studying factorizations also aids in understanding unitarity cuts at the loop level. Recently, factorizations have been utilized to determine Mellin amplitudes in AdS space (cf. Gonçalves et al. (2015); Alday et al. (2024); Cao et al. (2023)).
@@
Given the importance of factorizations in scattering amplitude studies, new types of factorizations have been investigated. In Cachazo et al. (2022), it was proposed that by setting specific Mandelstam variables to zero, amplitudes for biadjoint $\phi^{3}$ theory, the Nonlinear Sigma Model (NLSM), and a special Galileon theory Cachazo et al. (2015) split into three currents, a property termed semi-locality. This differs from conventional factorizations as no residues are computed. The smoothly splitting amplitudes for $\phi^{p}$ theories are discussed in Giménez Umbert and Yeats (2024).
@@
In Arkani-Hamed et al. (2023a), another form of three-part splitting was discovered for $\text{Tr}(\phi^{3})$ theory and NLSM. By setting certain two-point Mandelstam variables in a rectangle to zero and then turning back on one of them, the amplitudes split into three parts near their zeros. These two types of three-part splitting for scalar theories are unified by a so-called 2-split proposed in Cao et al. (2024a, b), which asserts that amplitudes factorize into two currents under certain subspaces of the Mandelstam variables. Applying the 2-splits in different ways reproduces both types of three-part splitting found in Cachazo et al. (2022) and Arkani-Hamed et al. (2023a). Additionally, significant progress has been made in extending such split factorizations to loop orders by analyzing the curve integral formulation for ${\rm Tr}(\phi^{3})$ amplitudes, as discussed in Arkani-Hamed et al. (2023a, b, 2024); Arkani-Hamed and Figueiredo (2024a, b).
@@
Beyond scalar theories, hidden zeros in tree-level Yang-Mills (YM) and GR amplitudes were also proposed in Arkani-Hamed et al. (2023a) by setting specific Mandelstam variables in a rectangle to zero along with their corresponding Lorentz products involving polarizations. Subsequently, the 2-split for YM and GR amplitudes was designed in Cao et al. (2024a, b) by turning on some vanishing Mandelstam variables while setting additional Lorentz products involving polarizations to zero. The hidden zeros in scalar theories have been further studied in Bartsch et al. (2024); Li et al. (2024); Rodina (2024) using double copy relations Kawai et al. (1986); Bern et al. (2008) and uniqueness bootstrap methods, which are also claimed to be applicable to the hidden zeros in amplitudes of YM, GR, and other theories.
@@
In this paper, we further investigate the new type of factorization of YM amplitudes by setting specific Mandelstam variables in a rectangle to zero, identical to the constraint used to produce zeros in certain scalar amplitudes Arkani-Hamed et al. (2023a). We found that under this constraint, YM amplitudes always factorize into sums of products of lower-point YM amplitudes. In addition to the summation over all possible states of the polarization of the internal gluon that is exchanged, as in conventional factorization, our factorization further includes a summation over different gluon pairs appearing in the vanishing Mandelstam variable rectangle. This new pattern of factorizations has been numerically verified for amplitudes up to and including 11 points.
@@
We refer to the contribution from each gluon pair as a gluon pair contribution. By eliminating all Lorentz products involving polarizations associated with the vanishing Mandelstam variables in the rectangle, we can observe how all gluon pair contributions vanish directly in our factorization formulas, thus making the hidden zeros manifest. Conversely, since the proposal of zeros in YM amplitudes in Arkani-Hamed et al. (2023a), an important question has been the behavior of amplitudes near these zeros when Lorentz products involving polarizations are turned on. Our factorization formulas address this question precisely, demonstrating how contributions accumulate as more Lorentz products are restored, ultimately recovering the full factorization formulas.
@@
Our factorization formulas for YM amplitudes differ from the 2-split proposed in Cao et al. (2024a, b) in that we only need to set certain Mandelstam variables to zero while allowing all Lorentz products involving polarization to remain. Additionally, our factorization formulas typically include a sum of terms instead of a single product of two amplitudes. However, combining our factorization formulas with the 2-split formulas yields another elegant result. By imposing conditions on Lorentz products such that only a single gluon pair contribution survives in our factorization summation for YM amplitudes, and then further eliminating additional Lorentz products involving polarizations, the entire YM amplitude factorizes into a sum of three currents. Continued study of our new factorization should uncover more properties of YM amplitudes, providing a novel description.
@@
We claim that the $n$-point YM amplitudes factorize into sums of products of a three-point and an $(n-1)$-point YM amplitude in the subspace of Mandelstam variables $h_{m}=0$ for $m=1,2,3,4$ and $n\geq m+3$,
@@
$F_{m,n}(i,j)=\frac{1}{s_{12\ldots m+1}}\sum_{\epsilon_{\hat{j}}}A(ij-\!\!\hat{j})\sum_{\rho\in S_{m-1}}X(s,\rho)$
@@
Here, we first sum over all gluon pairs $(i,j)$ appearing in the matrix $h_{m}$. Correspondingly, we call $F_{m,n}(i,j)$ as a gluon pair contribution. Then, we sum over all possible states of the polarization of the internal gluon $\hat{j}$ that is exchanged. The momentum conservation of the three-point amplitude $A(ij-\hat{j})$ implies that $k_{-\hat{j}}=-k_{\hat{j}}=k_{i}+k_{j}$, which is on-shell under the constraint $s_{ij}\in h_{m}=0$. Finally, we sum over all permutations $\rho$ of the $m-1$ gluons $1,2,\ldots,i-1,i+1,\ldots,m$.
The terms $X(s,\rho)$ are ratios of Mandelstam variables involving only the gluons $11,2,\ldots,m+$ 1. Their denominators consist of all planar poles between 1 and $m$ that include gluon $i$,
$X(s,\rho)=\frac{\tilde{X}(s,\rho)}{\prod_{\begin{subarray}{c}1\leq p\leq i\leq q\leq m\\ p<q\end{subarray}}s_{p\,p\!+\!1\ldots q}}\,.$ (4)
The numerators $\tilde{X}(s,\rho)$ depend on ordering $\rho$ and are polynomials of degree $(m-1)+(i-1)(m-i)$ in the Mandelstam variables $s_{rt}$ with $11\leq r<t\leq m+$ 1. To simplify their expressions, one may use the relations $s_{rn}=-\sum_{t=1,t\neq r}^{m+1}s_{rt}$ for $1\leq r\leq m$ on the support of $h_{m}=0$. Similarly, $s_{12\ldots m+1}=-s_{12\ldots mn}$ when $h_{m}=0$.
Here are some examples of $X(s,\rho)$ for $m=1,2$,
${\rm when}~{}m=1,i=1,\quad X(s,\emptyset)=-1\,,$
${\rm when}~{}m=2,i=1,\quad X(s,2)=-\frac{s_{12}+s_{23}}{s_{12}}\,,$
${\rm when}~{}m=2,i=2,\quad X(s,1)=-\frac{s_{12}+s_{1n}}{s_{12}}\,.$
We will present the expressions of $X(s,\rho)$ for $m=3,4$ later.
For $n\leq 11$, either the length $m$ or the width $m^{\prime}=n-m-2$ of the matrix $h_{m}$ in (1) is no more than 4. In cases where $m>4$ but $m^{\prime}\leq 4$, we can simply relabel $a\to a^{\prime}=n-a$ modulo $n$ for all $a=1,2,\ldots,n$ and apply our formulas (2) and (3) again to achieve the factorization of an amplitude under the constraint $h_{m}=0$. Many advanced methods have been developed to efficiently calculate high-point YM amplitudes (cf. Berends and Giele (1988); Mafra and Schlotterer (2016); Du and Teng (2017); Edison and Teng (2020); He et al. (2021); Arkani-Hamed et al. (2023b)). We have numerically verified our factorization formulas up to 10 points. By choosing specific numeric kinematic data, we further verified our formulas up to 11 points. Starting at $n=12$, there are cases where $m>4$ and $m^{\prime}>4$ simultaneously, necessitating the explicit expression of $X(s,\rho)$ for $m>4$, which we leave for future work.
@@
where the momentum $k_{\hat{j}}=k_{i}+k_{j}$ and polarization $\epsilon_{\hat{j}}=\epsilon_{j}$ of the new gluon $\hat{j}$ are on-shell and orthogonal to each other on the support of the conditions $H_{m}=0$ except $\epsilon_{i}\cdot k_{j}\neq 0$.
@@
In the following, we show more details for the cases $m=1,2,3,4$ respectively.
@@
$\quad-\frac{1}{s_{12}}\sum_{j=3}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(23\ldots j\!-\!1\,\hat{j}\,j\!+\!1\ldots n)\,.$
@@
$A(1234)$ $\xrightarrow{s_{13}=0}-\frac{1}{s_{12}}A_{\mu}(13;-\hat{3})A^{\mu}(42;\hat{3})\,,$
$A(12345)$ $\xrightarrow{s_{13}=s_{14}=0}-\frac{1}{s_{12}}\big{(}A_{\mu}(13;-\hat{3})A^{\mu}(452;\hat{3})$
$\qquad\qquad+A_{\mu}(14;-\hat{4})A^{\mu}(523;\hat{4})\big{)}\,.$ (12)
@@
$\quad-\frac{s_{12}+s_{23}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(234\ldots\hat{j}\ldots n)$
$\quad-\frac{s_{12}+s_{1n}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(2j\,-\!\hat{j})A(134\ldots\hat{j}\ldots n)\,,$
@@
$A(12345)$ $\xrightarrow{s_{14}=s_{24}=0}-\frac{s_{12}+s_{23}}{s_{12}s_{123}}A_{\mu}(14;-\hat{4})A^{\mu}(523;\hat{4})$
$\quad-\frac{s_{12}+s_{15}}{s_{12}s_{123}}A_{\mu}(24;-\hat{4})A^{\mu}(513;\hat{4}).$ (14)
@@
$+$ $A(3245\ldots{\hat{j}}\ldots n)s_{3n}s_{24}\Big{)}\,,$ (15)
$F_{3,n}(2,j)$ $=\frac{1}{s_{12}s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(2j-\!\hat{j})$
@@
-- $A(3145\ldots{\hat{j}}\ldots n)s_{14}s_{3n}(s_{12}+s_{23})\,\Big{)}\,,$ (16)
$F_{3,n}(3,j)$ $=\frac{1}{s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(3j-\!\hat{j})$
@@
$+$ $A(2145\ldots{\hat{j}}\ldots n)s_{14}s_{2n}\Big{)}\,.$ (17)
 Summing over all such contributions yields the factorization of YM amplitudes under the condition $h_{3}=0$ as shown in (2). The expressions for $X(s,\rho)$ for $m=3$ can be directly derived from these contributions. For instance, from the last line of (16), we obtain $X(s,31)=-s_{14}s_{3n}(s_{12}+s_{23})/(s_{12}s_{23}s_{123})$ for $F_{3,n}(2,j)$.
Note that $F_{3,n}(1,j)$ and $F_{3,n}(3,j^{\prime})$ with $j,j^{\prime}=5,6,\ldots,n-1$ and $j+j^{\prime}=n-4$ are identified by relabeling $a\leftrightarrow a^{\prime}=n+4-a$ modulo $n$ for all $1\leq a\leq n$. So are $F_{3,n}(2,j)$ and $F_{3,n}(2,j^{\prime})$. That is, $X(s,13)$ and $X(s,31)$, remain invariant under this relabeling up to an overall minus sign.
For $m=4$, each gluon pair contribution $F_{4,n}(i,j)$ involves six $(n{-}1)$-point amplitudes, $A(\rho(12\ldots i\!\!/\ldots 4)56\ldots\hat{j}\ldots n)$ with $\rho\in S_{3}$. Their explicit expressions are present in the supplemental materials.
@@
Although $s_{aj}\in h_{m}=0$, $s_{a\hat{j}}$ does not vanish, prompting the further factorization of the $(n-1)$-point amplitudes $A\big{(}\rho\ldots\hat{j}\ldots n\big{)}$ into a product of two currents based on the 2-split construction proposed in Cao et al. (2024a, b), with additional Lorentz products involving polarizations needing to be set to zero.
@@
Therefore, based on the 2-split construction proposed in Cao et al. (2024a, b), we just need to further impose the conditions
@@
$A(\mathbb{I}_{n})$ (18)${(H_{m}/\{x_{ij}\})=0}\frac{1}{s_{12\ldots m+1}}A^{\mu}(ij;-\hat{j})\Big{[}\sum_{\rho\in S_{m-1}}X(s,\rho)$ )
@@
Here, $J^{{\rm YM}+\phi^{3}}(n^{\phi}\ldots;\tilde{j}^{\prime}{}^{\phi})$ are single-trace currents for a mixed theory, dubbed as ${\rm YM}+\phi^{3}$, with two on-shell scalars $n,m+1$ and one off-shell scalar $\tilde{j}^{\prime}{}^{\phi}$. $J^{\rm YM}(j+1\ldots j-1;\tilde{j})$ is a YM current with an off-shell leg $\tilde{j}$. The momenta of the off-shell particles $\tilde{j}^{\prime}{}^{\phi}$ and $\tilde{j}$ are constrained by momentum conservation in their currents. In practice, one can use amplitude formulas to compute these currents, ensuring that all Mandelstam variables involving the off-shell particles and $s_{m\!+\!1\,n}$ are eliminated on the support of momentum conservation. For more details on the currents, refer to Cao et al. (2024a, b); Naculich (2015).

### v3

@@ Abstract
We propose a new factorization pattern for tree-level Yang-Mills (YM) amplitudes, where they decompose into a sum of gluings of two lower-point amplitudes by setting specific two-point non-planar Mandelstam variables within a rectangular configuration to zero. This approach manifests the hidden zeros of YM amplitudes recently identified. Furthermore, by setting specific Lorentz products involving polarization vectors to zero, the amplitudes further reduce to a sum of products of three currents. These novel factorizations provide a fresh perspective on the structure of YM amplitudes, potentially enhancing our understanding and calculation of these fundamental quantities.
@@ Introduction
At the loop level, (generalized) unitarity cuts are employed to determine the loop integrand of amplitudes Bern et al. (1994, 1995). The primary difference from conventional factorization is the substitution of a loop propagator for a tree propagator on-shell. Thus, studying factorizations also aids in understanding unitarity cuts at the loop level. Recently, factorizations have been utilized to determine Mellin amplitudes in AdS space (cf. Gonçalves et al. (2015); Alday et al. (2024); Cao et al. (2024a)).
@@
Given the importance of factorizations in scattering amplitude studies, new types of factorizations have been investigated. In Cachazo et al. (2022), it was proposed that by setting specific Mandelstam variables to zero, amplitudes for biadjoint $\phi^{3}$ theory, the Nonlinear Sigma Model (NLSM), and a special Galileon theory Cachazo et al. (2015) split into three currents, a property termed semi-locality. This differs from conventional factorizations as no residues are computed  [^1] The smoothly splitting amplitudes for $\phi^{p}$ theories are discussed in Giménez Umbert and Yeats (2024). See more recent progress in Giménez Umbert and Sturmfels (2025); Early (2025).
@@
In Arkani-Hamed et al. (2024a), another form of three-part splitting was discovered for $\text{Tr}(\phi^{3})$ theory and NLSM. By setting certain two-point Mandelstam variables in a rectangle to zero and then turning back on one of them, the amplitudes split into three parts near their zeros. These two types of three-part splitting for scalar theories are unified by a so-called 2-split proposed in Cao et al. (2024b, c), which asserts that amplitudes factorize into two currents under certain subspaces of the Mandelstam variables. Applying the 2-splits in different ways reproduces both types of three-part splitting found in Cachazo et al. (2022) and Arkani-Hamed et al. (2024a). Additionally, significant progress has been made in extending such split factorizations to loop orders by analyzing the curve integral formulation for ${\rm Tr}(\phi^{3})$ amplitudes, as discussed in Arkani-Hamed et al. (2024a, 2023a, b); Arkani-Hamed and Figueiredo (2024a, b).
@@
Beyond scalar theories, hidden zeros in tree-level Yang-Mills (YM) amplitudes were also proposed in Arkani-Hamed et al. (2024a) by setting specific Mandelstam variables in a rectangle to zero along with their corresponding Lorentz products involving polarizations. Subsequently, the 2-split for YM and GR amplitudes was designed in Cao et al. (2024b, c) by turning on some vanishing Mandelstam variables while setting additional Lorentz products involving polarizations to zero. The hidden zeros in scalar theories have been further studied in Bartsch et al. (2024); Li et al. (2024) using double copy relations Kawai et al. (1986); Bern et al. (2008) and uniqueness bootstrap methods Rodina (2025), which are also claimed to be applicable to the hidden zeros in amplitudes of YM, GR, and other theories.
@@
In this paper, we further investigate a new type of factorization of YM amplitudes by setting specific Mandelstam variables in a rectangle to zero, identical to the constraint used to produce zeros in certain scalar amplitudes Arkani-Hamed et al. (2024a). We found that under this constraint, YM amplitudes always factorize into sums of gluings of lower-point YM amplitudes. In addition to the summation over all possible states of the polarization of the internal gluon that is exchanged, as in conventional factorization, our factorization further includes a summation over different gluon pairs appearing in the vanishing Mandelstam variable rectangle.
[…]
Figure 1: New factorization of YM amplitudes: An $n$-point Yang-Mills amplitude factorizes into a sum over contributions, each involving a 3-point amplitude and an $(n-1)$-point amplitude. The sum runs over all permutations $\rho$ of the $(m-1)$ external gluons. The internal gluon $\hat{j}$ is exchanged in two diagrams. For simplicity, some rational functions of Mandelstam variables are omitted in this schematic representation.
The first graph in this figure also provides an intuitive explanation for the constraints on the Mandelstam variables in $h_{m}$, as defined in (1). Selecting any two non-adjacent external gluon legs (e.g., $m{+}1$ and $n$) in a color-ordered amplitude naturally partitions the remaining legs into two sets. Setting all two-particle Mandelstam variables that involve legs from both sets to zero gives $h_{m}=0$.
@@
We refer to the contribution from each gluon pair as a gluon pair contribution. By eliminating all Lorentz products involving polarizations associated with the vanishing Mandelstam variables in the rectangle, we can observe how all gluon pair contributions vanish directly in our factorization formulas, thus making the hidden zeros manifest. Conversely, since the proposal of zeros in YM amplitudes in Arkani-Hamed et al. (2024a), an important question has been the behavior of amplitudes near these zeros when Lorentz products involving polarizations are turned on back. Our factorization formulas address this question precisely, demonstrating how contributions accumulate as more Lorentz products are restored, ultimately recovering the full factorization formulas.
@@
Our factorization formulas for YM amplitudes differ from the 2-split proposed in Cao et al. (2024b, c) in that we only need to set certain Mandelstam variables to zero while allowing all Lorentz products involving polarization to remain. Additionally, our factorization formulas typically include a summation of gluing contributions instead of a single product of two amplitudes. However, combining our factorization formulas with the 2-split formulas yields another elegant result. By imposing conditions on Lorentz products such that only a single gluon pair contribution survives in our factorization summation for YM amplitudes, and then further eliminating additional Lorentz products involving polarizations, the entire YM amplitude factorizes into a sum of three currents. Continued study of our new factorization should uncover more properties of YM amplitudes, providing a novel description.
@@
We claim that the $n$-point YM amplitudes with canonical ordering factorize into sums of gluings of a three-point and an $(n-1)$-point YM amplitude in the subspace of Mandelstam variables $h_{m}=0$ for $m\geq 1$ and $n\geq m+3$,
@@
$F_{m,n}(i,j)=\frac{(-1)^{n+1}}{s_{12\ldots m+1}}\sum_{\epsilon_{\hat{j}}}A(ij-\!\!\hat{j})\sum_{\rho\in S_{m-1}}X(s,\rho)$
@@
This factorization, illustrated in fig. 1, involves summing over all gluon pairs $(i,j)$ appearing in the matrix $h_{m}$. Each term in the sum, denoted $F_{m,n}(i,j)$, represents a “gluon pair contribution”. Additionally, we sum over all possible polarization states of the internal gluon $\hat{j}$, which is exchanged between the two amplitudes.
Momentum conservation in the three-point amplitude $A(ij-\hat{j})$ ensures that the internal gluon satisfies $k_{\hat{j}}=-k_{-\hat{j}}=k_{i}{+}k_{j}$. Under the constraint $h_{m}=0$, the Mandelstam variable $s_{ij}$ vanishes, making the exchanged gluon on-shell. Finally, we sum over all permutations $\rho$ of the $m{-}1$ gluons $\{1,2,\dots,i{-}1,i{+}1,\dots,m\}$.
The terms $X(s,\rho)$ are ratios of Mandelstam variables that involve only the gluons $11,2,\ldots,m+$ 1 and they remain independent of the total number of particles $n$. Examples of $X(s,\rho)$ for low values of $m$ are as follows,
$\text{For }m=1,i=1,\quad X(s,\emptyset)=1\,,$
$\text{For }m=2,i=1,\quad X(s,2)=\frac{s_{12}+s_{23}}{s_{12}}\,,$
$\text{For }m=2,i=2,\quad X(s,1)=-\frac{s_{13}}{s_{12}}\,.$ (4)
In general, $X(s,\rho)$ is expressed as
[…]
×$\times\,{\cal B}_{m+1,i}[12\ldots i{-}1,i{+1}\ldots m|\pi]\,,$ (5)
where $g^{-1}$ is the inverse of the matrix $g[\rho,\pi]$, defined by
$g[\rho,\pi]:=$ $\sum_{\begin{subarray}{c}a=1\\ a\neq i\end{subarray}}^{m+1}\left(\sum_{b\in Y(a,\rho)}s_{ib}+s_{i\,m+1}\right)$
×$\times{\cal B}_{m+1,i}[Y(\rho,a),Y(a,\rho)|\pi]\,.$ (6)
[…]
The term ${\cal B}_{m+1,i}[\alpha,\beta|\pi]$ in (Main Claim) and (Main Claim) is derived from the coefficients in the BCJ relations for YM amplitudes, as outlined in Bern et al. (2024, 2008),
[…]
Here, $k_{\hat{n}}$, determined by momentum conservation, is typically on-shell to maintain gauge invariance in YM amplitudes. However, in this context, the on-shell condition for $k_{\hat{n}}$ is relaxed while ensuring that ${\cal B}_{m+1,i}[\alpha,\beta|\pi]$ remains independent of $k_{\hat{n}}$.
It turns out that the denominators of $X(s,\rho)$ correspond to all planar poles involving gluon $i$, ranging from $1$ to $m$. Numerically, this structure has been verified up to $m=7$. The numerators are polynomials in the Mandelstam variables $s_{rt}$ for $11\leq r<t\leq m+$ 1 and depend on the ordering $\rho$. Explicit forms of ${X}(s,\rho)$ with polynomial numerators for $m\leq 5$ are provided in ancillary files.
Finally, the contributions $F_{m,n}(i,j_{1})$ and $F_{m,n}(m+1-i,j_{2})$, where $j_{1},j_{2}=m+2,m+3,\ldots,n-1$ and $j_{1}+j_{2}=n+m+1$, are equivalent up to a sign under the relabeling $a\leftrightarrow m+1-a$. As a result, $X(s,\rho)$ satisfies the symmetry,
$X(s,\rho^{T}\big{|}_{a\to m+1-a\,{\rm mod}\,n})=X(s,\rho)\big{|}_{{a\to m+1-a\,{\rm mod}\,n}}\,,$ (8)
on the support of $h_{m}=0$, i.e. $s_{rn}=-\sum_{t=1,t\neq r}^{m+1}s_{rt}$ for $1\leq r\leq m$.
As an illustration, in (4), the term $X(s,1)$ can be rewritten as $\frac{s_{12}+s_{1n}}{s_{12}}$ on the support of $h_{2}=0$, making the symmetry (8) explicit when compared with the expression for $X(s,2)$.
A rigorous proof of the new factorization formula (2) is provided in the companion paper Zhang (2025).
@@
where the momentum $k_{\hat{j}}=k_{i}+k_{j}$ and polarization $\epsilon_{\hat{j}}=\epsilon_{j}$ of the new gluon $\hat{j}$ are on-shell and orthogonal to each other on the support of the conditions $H_{m}=0$ except $\epsilon_{i}\cdot k_{j}\neq 0$. If only $k_{i}\cdot\epsilon_{j}\in x_{ij}$ is turned on, $\epsilon_{\hat{j}}=-\epsilon_{i}$. When turning on $\epsilon_{i}\cdot\epsilon_{j}$ instead, $\epsilon_{\hat{j}}$ can take either $k_{i}$ or $-k_{j}$, reflecting the gauge redundancy of the amplitudes.
@@
In the following, we show more details for the cases $m=1,2,3,4,5$ respectively.
@@
$\quad\frac{(-1)^{n+1}}{s_{12}}\sum_{j=3}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(23\ldots j\!-\!1\,\hat{j}\,j\!+\!1\ldots n)\,.$
@@
$A(1234)$ $\xrightarrow{s_{13}=0}-\frac{1}{s_{12}}\sum_{\epsilon_{\hat{3}}}A(13-\!\hat{3})A(2\hat{3}4)\,,$
$A(12345)$ $\xrightarrow{s_{13}=s_{14}=0}\frac{1}{s_{12}}\big{(}\sum_{\epsilon_{\hat{3}}}A(13-\!\hat{3})A(2\hat{3}45)$
$\qquad\qquad+\sum_{\epsilon_{\hat{4}}}A(14-\!\hat{4})A(23\hat{4}5)\big{)}\,.$ (16)
@@
$\quad(-1)^{n+1}\frac{s_{12}+s_{23}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(234\ldots\hat{j}\ldots n)$
$\quad+(-1)^{n}\frac{s_{13}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(2j\,-\!\hat{j})A(134\ldots\hat{j}\ldots n)\,,$
@@
$A(12345)$ $\xrightarrow{s_{14}=s_{24}=0}\frac{s_{12}+s_{23}}{s_{12}s_{123}}\sum_{\epsilon_{\hat{4}}}A(14-\!\hat{4})A(23\hat{4}5)$
$\quad-\frac{s_{13}}{s_{12}s_{123}}\sum_{\epsilon_{\hat{4}}}A(24-\!\hat{4})A(13\hat{4}5).$ (18)
@@
$-A(3245\ldots{\hat{j}}\ldots n)s_{3n}s_{24}\Big{)}\,,$ (19)
$F_{3,n}(2,j)$ $=\frac{(-1)^{n+1}}{s_{12}s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(2j-\!\hat{j})$
@@
$+A(3145\ldots{\hat{j}}\ldots n)s_{14}s_{3n}(s_{12}+s_{23})\,\Big{)}\,,$ (20)
$F_{3,n}(3,j)$ $=\frac{(-1)^{n+1}}{s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(3j-\!\hat{j})$
@@
$-A(2145\ldots{\hat{j}}\ldots n)s_{14}s_{2n}\Big{)}\,.$ (21)
Summing over all such contributions gives the factorization of YM amplitudes under the condition $h_{3}=0$, as shown in (2). The explicit forms of $X(s,\rho)$ for $m=3$ can be directly derived from these contributions. For example, from the final term in (20), we find $X(s,31)=\frac{s_{14}s_{3n}(s_{12}+s_{23})}{s_{12}s_{23}s_{123}}$ for $F_{3,n}(2,j)$. Furthermore, it can be verified that $X(s,13)$ satisfies the symmetry property in (8) under the relabeling $a\leftrightarrow a^{\prime}=n+4-a\,{\rm mod}\,n$ for all $1\leq a\leq n$.
For $m=4$, each gluon pair contribution $F_{4,n}(i,j)$ involves six $(n{-}1)$-point amplitudes, $A(\rho(12\ldots i\!\!/\ldots 4)56\ldots\hat{j}\ldots n)$ with $\rho\in S_{3}$. Their explicit expressions are present in the supplemental materials. All ${X}(s,\rho)$ for $m\leq 5$ are also provided in ancillary files.
@@
Although $s_{aj}\in h_{m}=0$, $s_{a\hat{j}}$ does not vanish, prompting the further factorization of the $(n-1)$-point amplitudes $A\big{(}\rho\ldots\hat{j}\ldots n\big{)}$ into a product of two currents based on the 2-split construction proposed in Cao et al. (2024b, c), with additional Lorentz products involving polarizations needing to be set to zero.
@@
Therefore, based on the 2-split construction proposed in Cao et al. (2024b, c), we just need to further impose the conditions
@@
$A(\mathbb{I}_{n})$ (22)${(H_{m}/\{x_{ij}\})=0}\frac{(-1)^{n+1}}{s_{12\ldots m+1}}A^{\mu}(ij;-\hat{j})\Big{[}\sum_{\rho\in S_{m-1}}X(s,\rho)$ )
@@
Here, $J^{{\rm YM}+\phi^{3}}(n^{\phi}\ldots;\tilde{j}^{\prime}{}^{\phi})$ are single-trace currents for a mixed theory, dubbed as ${\rm YM}+\phi^{3}$, with two on-shell scalars $n,m+1$ and one off-shell scalar $\tilde{j}^{\prime}{}^{\phi}$. $J^{\rm YM}(j+1\ldots j-1;\tilde{j})$ is a YM current with an off-shell leg $\tilde{j}$. The momenta of the off-shell particles $\tilde{j}^{\prime}{}^{\phi}$ and $\tilde{j}$ are constrained by momentum conservation in their currents. In practice, one can use lower-point amplitude formulas to compute these currents, ensuring that all momenta of the off-shell particles including those in the Mandelstam variables and $s_{m\!+\!1\,n}$ are eliminated on the support of momentum conservation in each current. For more details on the currents, refer to Cao et al. (2024b, c); Naculich (2015).